OpenStudy (anonymous):

What is the probability of rolling a total of 7 in two rolls of a standard die if you get an even number on the first roll? Thanks~!

OpenStudy (zehanz):

To get 7, with an even number to begin with, look at the possibilites for this: 2+5, 4+3, 6+1. We already know that the first number is even, so the probability of it being 2, 4 or 6 is 1/3 each. The probability of getting 5, 3, or 1 is 1/6 each. Total probability: P(2,5)+P(4,3)+P(6,1)=(1/3)*(1/6)*3=1/6

OpenStudy (anonymous):

thank you!